### v2

@@ 4.3 Attack Experiment Design
We perform our attacks on an AI Chatbot designed to answer common medical queries from a patient. The user is a patient interacting with the chatbot and the attacker seeks to learn which query the user asks. We conduct the following three fingerprinting attacks:
@@
Experiment 1. The first experiment studies a scenario where the attacker has exact knowledge of the complete set of queries, allowing the attacker to profile the exact queries offline that will be leaked out in the online phase. In this, we use a set of 50 prompts from real world human-LLM interactions acquired from Han et al. (2023). These prompts ask a variety of questions about different common diseases (the list of prompts is provided in the Appendix).
Experiment 2. In this experiment, we choose 50 prompts with commonplace diseases from our dataset that all starts with “What are the symptoms of”. Therefore, all the prompts have a similar structure. We design this experiment to see if our attack can distinguish similar prompts.
The training and testing prompts of both Experiment 1 and 2 are the same. This illustrates the case where the attacker has exact knowledge of the set of predefined prompts.
@@
Experiment 3. The last experiment studies a scenario where the attacker has only approximate knowledge of the set of queries asked by the user. To emulate this, the attacker profiles queries that are semantically similar to the queries asked by the user (e.g., profiled query: “For how long should you take zolpidem?” and user: “How long is zolpidem typically prescribed for?”). We reuse the prompts from Experiment 1 as our training prompts, but for testing, i.e., the user prompts to be leaked, we rephrase the prompts in the training set using ChatGPT-4o by prompting it as "Can you rephrase the following 50 questions while keeping the diseases in each question the same?"
4.4 Methodology
@@
Target Models. We target three recent speculative decoding techniques: LADE Fu et al. (2024) that uses self-drafting for speculation, REST He et al. (2024) that uses retrieval from datastore for speculation, and BiLD Kim et al. (2023) which uses a smaller model for speculation. However, our results are as applicable to other techniques using speculation. We demonstrate results for LADE with TinyLlama 1.1B Chat V1.0 Zhang et al. (2024a), for REST with Vicuna 7B Chiang et al. (2023) as the target model and ShareGPT Aeala (2023) data-store containing 120K conversations used for speculation, and for BiLD with Llama2 7B Touvron et al. (2023) as the target model and TinyLlama 1.1B Zhang et al. (2024a) as the smaller model for speculation. We run LADE and REST on NVIDIA RTX4090 (24GB) and BiLD on NVIDIA A100 (40GB) GPUs with Question-Answer tasks, to simulate a chatbot under attack.
@@
Datasets and Classifier. For each experiment, we use 50 queries of varying complexity, length, and semantics (Appendix lists all the prompts). For our training set, in the offline profiling phase of the attack, we run each query 5 to 30 times and extract the packet size traces (5 to 30 traces per query) and train a Random Forest classifier using them. For the test set in the online phase, we use the corresponding 50 queries and 5 new traces per query, collecting 250 new data points for evaluation, and report accuracy and F1-scores. For our attack, we use scikit-learn’s Random Forest with 150 estimators, max depth of 15, minimum 10 samples per split, minimum 1 sample per leaf, and mean squared error as the loss function.
Accuracy (%) F1 Score
TPQ $\rightarrow$ 5 10 20 30 5 10 20 30
@@
LADE 53.6% 63.2% 65.5% 68% 0.49 0.59 0.63 0.66
REST 91.2% 97.6% 100% 99.6% 0.91 0.98 1.0 1.0
BiLD 28.4% 41.6% 49.2% 52% 0.27 0.43 0.52 0.55
@@
LADE 33.2% 53.2% 64.8% 70.8% 0.3 0.51 0.63 0.7
REST 96.4% 97.6% 100% 99.6% 0.96 0.98 1.0 1.0
BiLD 58.8% 68% 75.2% 77.6% 0.54 0.64 0.73 0.76
@@
LADE 15.2% 16% 13.6% 18.8% 0.13 0.13 0.1 0.16
REST 35.2% 38.4% 40% 39.6% 0.27 0.32 0.34 0.33
BiLD 13.2 % 20% 20.4% 20% 0.11 0.16 0.16 0.16
Table 1: Results of the Query Fingerprinting Attack on three speculative decoding techniques (LADE, REST, BiLD) with temperature of 0.8, with 5 to 30 Traces Per Query (TPQ) used for training the classifier.
4.5 Results
Exact Knowledge Scenarios (EK). In both Experiment 1 and 2, the attacker leaks the input prompt with high accuracy: with accuracy of up to 70.8%, 100%, and 77.6% for LADE, REST, and BiLD respectively, compared to random guessing (2% for a set of 50 prompts). In Experiment 1, where we study randomly selected prompts and the set of profiled and leaked prompts are identical, the attack success rates improve as the number of traces per query (TPQ) profiled increases, reaching up to 68% accuracy and an F1 score of 0.66 with 30 TPQ for LADE, reaching 99.6% accuracy with 30 TPQ and an F1 score of 1.0 for REST, and reaching 52% accuracy with 30 TPQ and an F1 score of 0.55 for BiLD. In Experiment 2, where all our prompts have a similar structure, the results are similar, with attack accuracy for LADE, REST, and BiLD reaching 70.8%, 99.6%, and 77.6% respectively. This indicates that our attack success rate is not influenced by the structure of the prompt (input), as long as the response (output) is different.
Approximate Knowledge Scenarios (AK). In Experiment 3, where the leaked prompts are not exactly the same as the profiled prompts, the attack success rates are lower (up to 18.8% for LADE, 20% BiLD, and 40% for REST), but still significantly better than random-guessing (2%). This demonstrates that models using speculative decoding are vulnerable to query fingerprinting attacks, even when the attacker just has approximate knowledge about the queries.
Overall, all of the attacks are more successful on REST than BiLD and LADE. This is because REST demonstrates higher correct speculation rates compared to BiLD and LADE, and therefore has more prompt-dependent variations in the tokens per iteration and the packet sizes, leading to the attack achieving higher accuracy on REST.
4.6 Ablation Studies
Table  2 presents an ablation study studying the attack accuracy as the temperature of our target model varies.
@@ 4.6 Ablation Studies
Table 2: Accuracy of the Query Fingerprinting Attack on LADE and REST at different temperatures (0.3, 0.6, 0.8, 1.0), with 30 Traces Per Query (TPQ) used to train the classifier.
@@
For LADE and BiLD, as the temperature increases from 0.3 to 1.0, the attack accuracy for both Experiment 1 and Experiment 2 decreases. This is because at lower temperatures, the generations and speculation patterns are more stable and alike. Therefore, the output tokens for the same prompt and the packet size traces are more similar across different runs, making the attack more successful.
For REST, the attack accuracy is relatively stable for both Experiment 1 and 2 despite the increasing temperature. This is likely because the features are so pronounced that it withstands the perturbations caused in the output by the increasing temperature.
@@
The attack accuracy in Experiment 3 has an interesting trend with temperature. Since the attacker only has approximate knowledge of the prompts (train and test set are different), the accuracy of this attack is much lower than Experiment 1 and 2. However, unlike Experiment 1 and 2, for Experiment 3 the attack accuracy has a moderate increase as the temperature of the model increases. This is because, at low temperatures, with limited variation in the training data, the classifier overfits on the training set, and at higher temperatures, this problem does not exist, allowing the attack to reach slightly higher accuracies.
5 Intellectual Property Leakage
Speculative decoding techniques may use carefully selected data or tuned hyper-parameters for achieving the highest correct speculation rates. In this section, we focus on how a malicious user can craft inputs to recover such data or hyper-parameters. We assume the malicious user can examine the output and count the number of tokens streamed back per iteration, as outlined in the threat model in Section 3-2.
5.1 Leaking Datastore Used for Speculation
Techniques like REST He et al. (2024) rely on retrieval from an external datastore to generate speculative tokens. These datastores can be populated with proprietary information to tailor model outputs to specific domains, or data collected from user interactions that is private. Patterns of correct speculations arising from such data can be exploited to leak the sensitive information in such datastores.
5.1.1 Background on REST’s Speculation
REST builds a datastore of phrases in the format of context and continuation pairs, where a preceding context and its following continuation are stored. During generation, REST retrieves the longest context matching with the tokens generated so far, and uses it to construct a speculative draft by selecting top prefixes from the retrieved continuations. This speculative draft is verified using a tree attention mechanism by the target model.
5.1.2 Attack Design
@@
We observe that any tokens that are correctly speculated, i.e., returned as a group of tokens per iteration have to exist in the datastore. We propose an attack where a malicious user crafts inputs and observes which tokens are correctly speculated and all such tokens can thus be leaked from the datastore. With this, the attacker can systematically extract sequences from the datastore.
We study the following attacker strategies for input generation to achieve high token leakage rates:
@@
2. Leveraging Common Words. The LLM is prompted to generate paragraphs containing the most frequently occurring words, taken from a dataset of top 10,000 English words Brants & Franz (2021). This is based on the insight that there is a high chance of phrases in the datastore containing these words.
@@
3. Reusing Leaked Sequences. The LLM is prompted to generate paragraphs that extend phrases leaked out so far, as shown in Figure 5. This approach is based on the insight that the sequence that is already leaked out may be a part of a longer phrase in the datastore. Therefore, with such iterative prompting, additional tokens from the phrase can be leaked out in a focused manner, leading to more unique tokens being leaked with a similar number of queries.
Figure 5: Leakage process where an attacker exploits leaked sequences to extract longer datastore sequences.
Figure 6: Unique sequences leaked from REST datastore as time increases with different attack strategies, averaged over 3 runs per strategy.
5.1.3 Results
We test our attacks on a REST implementation using the ShareGPT datastore of 120K conversations for speculation on NVIDIA RTX 4090. Figure 6 shows the unique sequences leaded by each of the three attack strategies, averaged over three runs per strategy.
Random generation leaks 20,000 unique sequences in 20 minutes, but subsequently the rate slows down leaking only 70,000 sequences in 3 hours. In comparison, prompting with frequently occurring words leaks 31,000 unique sequences in 20 minutes, which increases linearly to 190,000 after 3 hours. Lastly, prompting with leaked sequences results in the most efficient leakage, leaking about 35,000 unique sequences in 20 minutes and about 200,000 after 3 hours, highlighting the amplification potential of feedback. These results show that an attacker can easily leak REST’s datastore when they can observe correct speculation patterns.
5.2 Leakage of Speculation Hyper-parameters
5.2.1 Background on LADE’s Implementation
5.2.2 Leaking Hyper-parameter N in LADE
As each iteration outputs at most N-1 tokens under correct speculation, we expect the maximum number of tokens per iteration to be N-1. Therefore, we prompt LADE with a simple prompt, “Repeat Letter ‘A’ 60 times” that should sustain the maximum correctly speculated tokens per iteration. We observe the maximum number of tokens per iteration with this prompt to deduce N-1, and thus learn N.
Figure 7 shows the number of correct speculated tokens for LADE with three different values of $N$ for this prompt. We see that for each value of N (4, 5, and 6), this prompt has a maximum of 3, 4, and 5 tokens per iteration (N-1), letting us learn the value of N as 4, 5, or 6 in each of these cases. This technique can be extended to learn any value of N used in LADE.
Figure 7: Number of tokens per iteration with LADE for different N, with a prompt that achieves maximum correct speculation (“Repeat Letter ‘A’ 60 times”. The maximum token count per iteration corresponds to value of N-1.)
5.2.3 Leaking Hyper-parameter G in LADE
To leak G, the number of candidate predictions considered for speculation, we force LADE to incur targeted mis-speculations. Here, we prompt the model to repeat a sequence of phrases with the same prefix (“I run”), i.e., “I run to the grocery shop. I run as fast as cars. I run with my best friend. I run from the lecture hall.". If the number of such phrases is less than or equal to G, LADE caches all the N-grams following the key “run”, resulting in it being able to correctly speculate the token that follows “run” in each of these phrases. However, if the number of phrases are greater than G, then with LRU replacement, newer n-grams evict older n-grams for the key “run” from the set of candidates which has a limited capacity of G. This causes the token following “run” to always be mis-speculated.
Figure 8 shows the traces of token generation time for LADE ($G=3$), as the number of phrases ($P$) of the type “I run …” increases from 2 to 4. When $P$ is 2 or 3 ($\leq G$), we notice all the tokens after the initial ones are correctly predicted in the generation (high number of tokens per iteration). But when $P$ is 4 ($>G$), every 7th token, i.e. the token following “run” is mis-speculated (one token per iteration). Thus, we leak the value of G to be 3. This attack can be generalized to any value of parameter G in LADE.
Figure 8: Time Per Token vs Token-ID for LADE with Guess Set (G = 3), using a prompt that asks the LLM to repeat P phrases sharing a common prefix “I run”. Periodic mis-speculations occur for the token after “run”, only when $P>G$, leaking the value of $G$.

### v3

@@ 4.3 Attack Design
We attack an AI Chatbot answering medical queries from a patient. The attacker seeks to learn which query a patient asks from a finite set of queries the model can answer, to potentially leak a disease the patient has, a privacy violation by law in geographies like EU or USA. We perform three attacks:
@@
Experiment 1. The attacker has exact knowledge of the complete set of queries, allowing the attacker to profile the exact queries offline and leak them out in the online phase. We choose a set of 50 prompts from a real-world human-LLM interactions dataset Han et al. (2023). These prompts ask a variety of questions about different diseases (the list of prompts is in Appendix A).
Experiment 2. Next, we ask the question, do the speculation patterns depend on structural patterns or deeper semantic meanings within the response? To analyze this better, we choose 50 prompts about diseases that all starts with “What are the symptoms of” from the same dataset, ensuring all prompts have a similar structure. With this, we seek to understand whether our attack can be used to leak deeper semantic meanings associated with the prompts (e.g., diseases of the prompts) even when all the prompts have similar structures.
@@
Experiment 3. Assuming that speculation patterns rely on semantic meanings of responses, can the attacker use this to leak private information about the prompt (e.g., user’s disease) without the exact knowledge of the user’s query? While Experiment 1 and 2 assumed training and testing prompts to be the identical, here in Experiment 3, the attacker only has approximate knowledge of user queries during profiling. The attacker’s training set consists of queries that are semantically similar to user’s test-set queries (e.g., profiled query: “For how long should you take zolpidem?” and user’s query: “How long is zolpidem typically prescribed for?”). We use the prompts from Experiment 1 as our training prompts, but during testing, we rephrase the prompts using ChatGPT-4o by prompting it as ”Can you rephrase the following 50 questions while keeping the content in the question the same?”. We provide a more general experiment with an out-of-distribution training set in Section 4.8.
4.4 Experimental Setup
@@
Target Models. We target four representative speculative decoding techniques: LADE Fu et al. (2024) and EAGLE Li et al. (2024a) using self-drafting, REST He et al. (2024) using retrieval and tree-verification, and BiLD Kim et al. (2023) using a smaller draft model. We run LADE with TinyLlama 1.1B Chat V1.0 Zhang et al. (2024a), REST with Vicuna 7B Chiang et al. (2023) as the target model and ShareGPT Aeala (2023) data-store of 120K conversations, and BiLD with Llama2 7B Touvron et al. (2023) and TinyLlama 1.1B Zhang et al. (2024a) as the target and draft model respectively. We run LADE and REST on NVIDIA RTX4090 (24GB) and BiLD on NVIDIA A100 (40GB) GPUs with Question-Answer tasks, to model a chatbot under attack.
@@
Datasets and Classifier. For each experiment, we use 50 queries of varying complexity, length, and semantics (Appendix A lists the prompts). For the training set in the profiling phase of the attack, we run each query 5 to 30 times (with varying temperatures) and extract the packet size trace (5 to 30 traces per query) and train a Random Forest classifier. For the test set in the online phase, we use the corresponding 50 queries and 5 new traces per query, collecting 250 data points for evaluation, and report accuracy. For our attack, we use scikit-learn’s Random Forest with 150 estimators, max depth of 15, minimum 10 samples per split and 1 sample per leaf, and mean squared error as loss function.
Traces Per Query (TPQ)
5 10 20 30
@@
LADE 53.6% 63.2% 65.5% 68%
REST 91.2% 97.6% 100% 99.6%
BiLD 28.4% 41.6% 49.2% 52%
@@
LADE 33.2% 53.2% 64.8% 70.8%
REST 96.4% 97.6% 100% 99.6%
BiLD 58.8% 68% 75.2% 77.6%
@@
LADE 15.2% 16% 13.6% 18.8%
REST 35.2% 38.4% 40% 39.6%
BiLD 13.2% 20% 20.4% 20%
Table 1: Accuracy of the Query Fingerprinting Attack on LADE, REST, and BiLD with temperature of 0.8, using 5 to 30 Traces Per Query (TPQ) for training.
@@
Table 2: Ablation Study. Attack accuracy of query fingerprinting as temperatures vary (0.3, 0.6, 0.8, 1.0), using 30 Traces Per Query (TPQ) for training.
4.5 Results
Exact Knowledge Scenarios. In both Experiment 1 and 2, the attacker leaks the input prompt with high accuracy: with accuracy of up to 70.8%, 100%, and 77.6% for LADE, REST, and BiLD respectively, compared to random guessing (2% for a set of 50 prompts). In Experiment 1, where we study randomly selected prompts and the set of profiled and leaked prompts are identical, the attack success rates improve as the number of traces per query (TPQ) profiled increases, reaching up to 68% accuracy and an F1 score of 0.66 with 30 TPQ for LADE, reaching 99.6% accuracy with 30 TPQ and an F1 score of 1.0 for REST, and reaching 52% accuracy with 30 TPQ and an F1 score of 0.55 for BiLD. In Experiment 2, where all our prompts have a similar structure but different semantic meaning, the results are similar, with attack accuracy for LADE, REST, and BiLD reaching 70.8%, 99.6%, and 77.6% respectively. This indicates that our attack success rate is not influenced by the structure of the prompt, but relies and leaks the semantic meanings of the prompt and response.
Approximate Knowledge Scenario. Based on the above results, Experiment 3 uses profiled prompts that are not identical to the leaked prompts, but still semantically similar. Here, the attack success rates are lower (up to 18.8% for LADE, 20% BiLD, and 40% for REST), but still significantly better than random-guessing (2%). Thus, speculative decoding can still leak the topic of a user’s prompt (e.g., user’s disease), with queries that are approximately similar in semantics to the user queries. Section 4.8 details a broader out-of-distribution experiment, where the training set is taken from public sources independent of the test set.
The attacks are more successful on REST than BiLD and LADE. This is because REST has higher correct speculation rates and more stable speculation patterns (due to its reliance on retrieval from a datastore of previous conversations) compared to BiLD and LADE, enabling higher attack accuracies.
4.6 Ablation Studies
@@
Table  2 presents an ablation study for the attack accuracy as the temperature of our target model varies. For LADE and BiLD, as the temperature increases from 0.3 to 1.0, the attack accuracy for both Experiment 1 and 2 decreases. This is because at lower temperatures, the generations and speculation patterns are more stable and alike. Hence, the output tokens for the same prompt and the packet sizes are more similar across runs, making the attack more successful. For REST, the attack accuracy is stable for both Experiment 1 and 2 despite the increasing temperature. This is because the features are so pronounced that it withstands perturbations in the output caused by the increasing temperature.
@@
In Experiment 3, the attacker accuracy is lower than Exp 1 or 2, as it only has approximate knowledge of the prompts (train and test set are different) However, unlike Experiment 1 and 2, for Experiment 3 the attack accuracy increases moderately as the temperature increases. This is because, at low temperatures, with limited variation in the training data, the classifier overfits on the training set, and at higher temperatures, this problem does not exist, allowing the attack be slightly more accurate.
4.7 Attack on Remote vLLM Server
To show real-world practicality, we demonstrate our attack on a remote LLM inference server on vLLM, with Llama3 8B Instruct as the target model and EAGLE speculative decoding. Our LLM server is hosted on a cloud-based A100 GPU, while our client is located in an academic network more than 1500 miles away. Our attacker snoops the encrypted network packets in between the server and the client, using Wireshark, and captures the bytes per packet. We also compare our attack with prior work, Carlini & Nasr (2024), that used packet inter-arrival-times as a side-channel, that varies based on inference-time optimizations. To mimic their setup, which focused on ChatGPT, that sends tokens generated in the same iteration in separate packets, we modify vLLM to similarly send separate packets for individual tokens generated in the same iteration. For our packet size side channel, we generate our trace by combining sizes of packets received within 55 ms of each other (based on observed network latencies).
Table 3 shows that, under no additional server load, our attack achieves a maximum of 77.6% accuracy, significantly outperforming prior work which achieves a maximum accuracy of 14.4%. This is because inter-arrival times provide a noisy signal due to varying network contentions. Moreover, under spikes in server load, the inter-arrival times (time per output token) can significantly spike as observed in recent work Zhang et al. (2024b), causing test-time observations to vary dramatically compared to training time. Consequently, assuming a standard log-normal distribution of time-per-iteration to mimic high server load, the accuracy of prior timing channel attacks drops to 6.4%. In comparison, our attack accuracy is unaffected, since we rely on the packet size side-channel, which remains robust to variations in server loads or network congestion.
No Server Load High Server Load
Temperature $\rightarrow$ 0.3 0.6 0.8 1.0 0.3 0.6 0.8 1.0
Packet Size Side Channel (ours) 77.6% 65.6% 44.8% 24.0% 77.6% 65.6% 44.8% 24.0%
Inter-Arrival Time Side Channel 14.4% 6.0% 5.2% 4.8% 6.4% 4% 2.4% 2.8%
(by Carlini & Nasr (2024))
Table 3: Accuracy of query fingerprinting attack (experiment 1) on vLLM with EAGLE’s speculative decoding at different temperatures (0.3, 0.6, 0.8, 1.0), using 30 Traces Per Query (TPQ) for training. We evaluate the attack using our packet-size side channel, and with inter-arrival times based side channel from prior work Carlini & Nasr (2024).
4.8 Out-of-Distribution Training
To study the effect of distribution mismatch between profiling and attack phases, we conducted an additional experiment where the attacker trains on an out-of-distribution (OOD) set of queries. Shown in  A.5, we used GPT-4o to generate 50 common diseases that users typically ask AI chatbots about, prepending the phrase “what are the symptoms of” to each. This OOD set was used for training, while evaluation was performed on the same 50 disease-related queries from Experiment 1 (with model temperature of 0.3 to 0.8), ensuring no prior knowledge of the test dataset during training.
Despite the lack of prior knowledge of the true query set, the attack substantially outperformed random guessing ($\sim$2%). LADE achieved 23% - 25% accuracy, REST 35% - 36%, and BiLD 23% - 25%, where success was defined as predicting a disease with similar symptoms (as per GPT-5 mini) compared to the ground truth, as temperature of the target model varies between 0.3 and 0.8 (see Table  7 in Appendix E for detailed results). These results show that although accuracy drops relative to the in-distribution setting, speculative decoding still leaks strong and learnable signals even under OOD conditions, without being much affected by temperature, underscoring the robustness of our side-channel attack.
5 Leaking Private Data Used for Speculation
Speculative decoding techniques may use carefully tuned data to ensure high correct speculation rates, which may contain private user data or other intellectual property. In this section, we show how a malicious user can leak out private data used for predictions. We show this on REST He et al. (2024), which relies on retrieval from a datastore to generate speculative tokens. This datastore can be populated with proprietary information to tailor model output to specific domains, or data collected from user interactions, which may be leaked by our attack. A similar attack can also leak speculation hyperparameters, which can be a company’s intellectual property, as shown in Appendix C.
@@ 5 Leaking Private Data Used for Speculation
Attack Design. REST retrieves the longest matching sequence from its datastore based on the tokens generated so far and uses the subsequent continuations as speculative tokens. Any tokens that are correctly speculated, i.e., returned as a group of tokens per iteration, have to exist in the datastore. We propose an attack where a malicious user crafts inputs and observes which tokens are correctly speculated and systematically leaks sequences of tokens from the datastore. We study the following attacker strategies for input generation to achieve high token leakage rates:
@@
2. Leveraging Common Words. The LLM is prompted to generate paragraphs containing the most frequently occurring words, drawn from a dataset of the 10,000 most common English words Brants & Franz (2021). This exploits the high likelihood that phrases in the datastore contain these words.
@@
3. Reusing Leaked Sequences. The LLM is iteratively prompted to generate text extending phrases already leaked out, as shown in Figure 3(a). This approach is based on the insight that the sequence that is already leaked out may be a part of a longer phrase in the datastore.
(a) Exploiting leaked sequences for longer extraction.
(b) Unique Sequences Leaked vs Time (avg of 3 runs).
Figure 3: Analysis of datastore leakage from REST.
Results. We test our attacks on REST using the ShareGPT datastore of 120K conversations for retrieval on NVIDIA RTX 4090. Figure 3(b) shows the unique sequences leaked by each attack strategy, averaged over three runs. Random generation initially leaks 20K sequences in 20 minutes, but slows down, leaking only 70K sequences in 3 hours. Leveraging common words improves this to 31K unique sequences leaked in 20 minutes and 190K after 3 hours. Lastly, prompting with leaked sequences is the most effective, leaking 35K unique sequences in 20 minutes and 200K in 3 hours, demonstrating how feedback amplifies leakage. These results show that a malicious user can reliably extract REST’s datastore via speculation patterns.